OpenStudy (anonymous):

Thought I was finally getting the hang of these... What am I doing wrong? (trig integral)

hartnn (hartnn):

question?

OpenStudy (anonymous):

yeah... I'll upload an image :p

hartnn (hartnn):

everything seems to be correct! just add +c in the end

hartnn (hartnn):

why do u feel u are wrong?

hartnn (hartnn):

oh, wait sec^2 x = 1+ tan^2 x

OpenStudy (anonymous):

ohhhhhhh.... I think that's my problem.... I messed up my trig identities. What is the correct identity?

hartnn (hartnn):

so there will be + sign in between 1+tan^2 x = sec^2 x

OpenStudy (anonymous):

ah ok :). Thank you. I need to commit these to memory... I'm not too good with trig identities... Also, there's one integral that I haven't tried yet that is int_sin(3x)cos(5x)... Is there some kind of identity to solve that integral?

hartnn (hartnn):

sin A cos B = ?

hartnn (hartnn):

sin A cos B = 0.5*(sin(A+B)+sin (A-B)) now can u solve ?

OpenStudy (anonymous):

Will that work even if the two quantities inside the trig functions are different?

hartnn (hartnn):

A and b are different...

OpenStudy (anonymous):

ok thanks :). I'm about to do that problem, so I will see how it goes :P. Thank you for the help too. I've lost count of how many times you've helped me haha.

hartnn (hartnn):

welcome :) no need to keep count.